# BUFFER-X: Towards Zero-Shot Point Cloud Registration in Diverse Scenes

### v1

@@ 5.4 Limitations
Table 4: Ablation study: the impact of different scale combinations on registration performance in the 3DMatch [88].
Fig. 8: Area plot of computation time per scale for each module on an NVIDIA GeForce RTX 3090 using 3DMatch [88]. Note that our cross-scale consensus maximization (red) only took 0.40 msec on average.
@@ A Parameters setup and pseudo code
1 Input:  Source cloud $\mathcal{P}$ and target cloud $\mathcal{Q}$; User-defined parameters $\tau_{v}$, $\delta_{v}$, $\delta_{r}$, $[\tau_{l},\tau_{m},\tau_{g}]$, and $N_{\text{FPS}}$
@@ 3 $\mathcal{P}_{r}\leftarrow\texttt{select\_larger\_cloud}(\mathcal{P},\mathcal{Q})$
4 $\mathcal{P}_{\text{sampled}}=\texttt{sample}(\mathcal{P}_{r},\delta_{v})$ Sample $\delta_{v}$of cloud points
@@
6 $v=\texttt{calc\_voxel\_size}(\mathcal{P}_{\text{sampled}},\tau_{v})$ See Eq. 3
7 $\mathcal{P}\leftarrow f_{v}(\mathcal{P}),\mathcal{Q}\leftarrow f_{v}(\mathcal{Q})$ Downsample the point clouds
@@ 8 $\mathcal{P}_{r}\leftarrow\texttt{select\_larger\_cloud}(\mathcal{P},\mathcal{Q})$
12 $\mathcal{M}^{\mathcal{P}}=\varnothing,\mathcal{M}^{\mathcal{Q}}=\varnothing$ Containers of embedding output
@@ 23 $\mathcal{D}=\varnothing,\mathcal{T}=\varnothing$
36 $\mathcal{I}=\texttt{consensus\_maximization}(\mathcal{D},\mathcal{T})$ See Eq. 9
@@
B Detailed experimental setups and datasets
Param. Description Value
$\tau_{v}$ Threshold in Eq. 3 0.05
$\tau_{l}$ Threshold for the local ($l$) search radius 0.005
$\tau_{m}$ Threshold for the middle ($m$) search radius 0.02
$\tau_{g}$ Threshold for the local ($g$) search radius 0.05
$\delta_{v}$ Sampling ratio for sphericity-based voxelization 10%
$N_{r}$ Number of sampling points for radius estimation 2,000
$r_{\text{max}}$ Maximum radius 5.0 m
$N_{\text{FPS}}$ Number of sampled points by FPS 1,500
$N_{\text{patch}}$ Maximum number of points in each patch 512
$\delta$ Truncation threshold for Huber loss 1.0
$H$ Height of the cylindrical map in Mini-SpinNet 7
$W$ Sector size of the cylindrical map in Mini-SpinNet 20
$D$ Feature dimension of the cylindrical map in Mini-SpinNet 32
  • 3DMatch follows conventional protocol proposed by Zeng et al.[88].
  • 3DLoMatch follows conventional protocol proposed by Huang et al.[32]. This dataset is derived from 3DMatch by selectively extracting pairs with low overlap (i.e., 10–30%), allowing for the evaluation of robustness to low-overlap scenarios.
  • ScanNet++F is from ScanNet++ [81]. For each sequence, a pre-merged PLY file representing the entire space is provided, along with the poses of the stationary FARO LiDAR sensor used to measure it. That is, the dataset does not provide individual scans. For that reason, using the given scanner poses, we generate per-scanner point clouds by sampling along raycasting paths from each position within the full map, simulating virtual scans. The sampling process respects the scanner’s horizontal angular resolution as well as its vertical resolution between consecutive rays [64], ensuring a realistic approximation of the original scans. For each ray, the nearest intersecting point in the merged point cloud is selected to simulate the original scan.
  • ScanNet++i is also from ScanNet++ [81]. There are depth images captured using a LiDAR sensor attached to an iPhone 13 Pro. These depth images were converted into point clouds using the toolbox provided by 3DMatch [88]. To generate dense point cloud fragments, 50 consecutive frames were accumulated. Finally, pairs with an overlap ratio of at least 0.4 (i.e., 40% overlap between two fragments) were selected as the final test pairs.
  • TIERS consists of Indoor06, Indoor08, Indoor09, Indoor10, and Indoor11 sequences in the TIERS dataset [60]. To reduce redundant test pairs, we exclude Indoor07 because it was acquired in the same room as Indoor06.
@@ B Detailed experimental setups and datasets
    Dataset/Sequence Name 3DMatch, 3DLoMatch ScanNet++F ScanNet++i TIERS
@@
    Ouster0-128
    Approx. Range [m] 3.5 7.0 3.5 110
    # of Points/Frame 336,274 1,381,013 444,876 34,777
    # of Test Pairs 1,623 / 1,781 2,016 2,074 870
@@
    Follow Zeng et al. [88]’s
@@
    Employed Sensor(s) Laser Bear Honeycomb Velodyne HDL-64E Rotating Hokuyo UTM-30LX
    Livox Avia,
    Aeva Aeries II,
@@
    Table A2: Overview of the datasets used in our experiments, including their environments, acquisition sites, measurement types, employed sensors, approximate ranges or scales, and the number of test pairs. The datasets cover a variety of indoor and outdoor environments, spanning different geographic and cultural regions.
  • WOD follows protocol proposed by Liu et al.[44]. This dataset is from the Waymo Open Dataset[69] Perception dataset by extracting LiDAR sequences and corresponding pose files, which are then converted into the KITTI format to ensure compatibility with standard benchmarking pipelines. We set $\tau_{\text{dist}}=10\,\text{m}$
  • KITTI follows conventional protocol proposed by Yew et al.[82]. In test scenes, 08, 09, and 10 sequences are employed. Originally, $\tau_{\text{dist}}=10\,\text{m}$.
  • ETH is from the gazebo_summer, gazebo_winter, wood_autmn, and wood_summer sequences of the dataset proposed by Pomerleau et al. [56]. The original dataset contains a wider various of scenes; however, following the existing protocol proposed by Ao et al. [5], we used four sequences.
  • KAIST is from the KAIST05 sequence of the HeLiPR dataset [34]. Originally, HeLiPR contains multiple sequences, but each sequence in the HeLiPR is much longer than those in MIT and Oxford, resulting in many more test pairs compared to other campus scenes (i.e., 1991 vs. 230 or 301 in Table A2). For this reason, we balanced the datasets by using only one sequence. We set $\tau_{\text{dist}}=10\,\text{m}$.
    Instead of emphasizing that HeLiPR is a heterogeneous LiDAR sensor dataset, we refer to a subset of it as KAIST in our paper to highlight that our dataset was curated with consideration for geographic and cultural environments. Similarly, we use MIT (from the Kimera-Multi dataset [71]) and Oxford (from the NewerCollege dataset [61]) to emphasize the institutions for the same reason. Further details are provided in Appendix C.
  • MIT is from 10_14_acl_jackal sequence of from the Kimera-Multi dataset [71], which is a multi-robot multi-session SLAM dataset. We could have used more scenes, but we chose to use only one sequence to a) match Oxford’s frame count as closely as possible (i.e., 230 vs. 301 in Table A2) and b) reduce the redundant test pairs, as multi-robot SLAM datasets often observe the same space multiple times.
@@
  • Oxford is from the 01, 05, and 07 sequences of the NewerCollege dataset [61]. An interesting aspect is that 01 and 07 were acquired by handheld setup, while the 05 sequence was acquired using a quadruped robot. As shown in Fig. A2, the campus scale was relatively small, so if we use only a single sequence, it only generates few test pairs. For that reason, to ensure at least a similar number of test pairs as MIT, unlike KAIST and MIT, we used three sequences and set $\tau_{\text{dist}}=5\,\text{m}$.
@@ C Rationale behind dataset selection
Fig. A2: Scale comparison of three sequences: Oxford from the NewerCollege [61], MIT from the Kimera-Multi [71], and KAIST from the HeLiPR dataset [34] at the same scale (i.e., 500 m). Note that although these sequences fall under the same campus category, their scales differ. For clarity, the map clouds are visualized with respect to their $z$ values.
@@
Acquisition setups.   We also considered that acquisition setups vary in multiple ways. A common bias in previous evaluations in the state-of-the-art approaches is that indoor scanning is performed using handheld devices, whereas outdoor scanning is conducted using vehicles (i.e., the assumption that indoor scanning is performed using a handheld setup, while outdoor scanning is conducted using a vehicle). Thus, we aimed to evaluate whether registration remains robust to different acquisition setups to challenge this assumption. For this reason, we included TIERS, which was acquired using a sensor cart, and Oxford, which was acquired using both a handheld device and a quadruped robot.
Additionally, MIT was captured using a mobile robot, which performs planar motion similar to a vehicle. However, due to its smaller size, the robot’s body experiences significantly more roll and pitch motion, potentially introducing greater motion noise compared to a vehicle.
Fig. A3: Examples of visualized LiDAR scans from different LiDAR sensors in (a) the Indoor10 of TIERS dataset [60] and (b) KAIST05 sequence of the HeLiPR dataset [34]. Note that even in the same environment, differences in the number of LiDAR rays and field of view result in point clouds with different patterns.
Diversity of geographic and cultural environments. Lastly, we also aimed to assess whether the acquired data could be generalized across geographic and cultural differences. Some studies [13] have claimed that training on KITTI and testing on KITTI-360 could demonstrate domain generalization. However, since both datasets were collected in Germany using the same vehicle platform, they fall short of demonstrating actual cultural differences.
@@ D Trade-off between generalization and robustness against partial overlaps
Here, as a supplement to the explanation in Sec. 5.4, we explain with an example why our methodology performed relatively poorly on 3DLoMatch in detail [32]. Suppose we have two arbitrary L-shaped point clouds, as presented in Fig. A4(a). Then, these two L-shaped point clouds can be registered in the following cases:
@@
In addition, even with a learnable non-linear robust kernel, this problem cannot be perfectly resolved. For instance, we examine two renowned non-convex cost functions: a) German-McClure (GM) function and b) truncated least squares (TLS) function, and use them as surrogate cost functions $\rho_{\mu}(r)$ that adjust their non-linearity by changing the control parameter $\mu$ [75]. Formally, by letting the $r$ be the residual and $\bar{c}$ be the user-defined parameter that determines the shape of a kernel, the GM function with $\mu$ can be expressed as follows:
@@
E Qualitative results in diverse scenes
We present the quantitative results in Fig. A5 and Fig. A6. Remarkably, even when trained on 3DMatch, which consists solely of RGB-D point clouds with an approximate maximum range of 3.5 m, the model performs robustly on sparse LiDAR point clouds in both indoor and outdoor environments. In particular, our approach successfully performed registration regardless of the sensor type or acquisition setup, whether the sparse 3D point clouds were acquired by solid-state LiDAR sensors (the first row among the KAIST rows in Fig. A6) or captured from a robot platform (MIT rows in Fig. A6).

### v3

@@ 5.4 Limitations
Table 3: Ablation study: the impact of different scale combinations on registration performance in the 3DMatch [Zeng17cvpr-3dmatch].
Fig. 8: Area plot of computation time per scale for each module on an NVIDIA GeForce RTX 3090 using 3DMatch [Zeng17cvpr-3dmatch].
@@ 6 Conclusion
Acknowledgments
This work was supported by IITP grant (RS-2021-II211343: AI Graduate School Program at Seoul National University) (5%), and by NRF grant funded by the Korea government (MSIT) (No. 2023R1A1C200781211 (65%) and No. RS-2024-00461409 (30%), respectively).
Appendix
@@ A Parameters setup and pseudo code
B Detailed experimental setups and datasets
  • 3DMatch follows conventional protocol proposed by Zeng et al.[Zeng17cvpr-3dmatch].
  • 3DLoMatch follows conventional protocol proposed by Huang et al.[Huang21cvpr-PREDATORRegistration]. This dataset is derived from 3DMatch by selectively extracting pairs with low overlap (i.e., 10–30%), allowing for the evaluation of robustness to low-overlap scenarios.
  • ScanNet++F is from ScanNet++ [Yeshwanth23iccv-Scannet++]. For each sequence, a pre-merged PLY file representing the entire space is provided, along with the poses of the stationary FARO LiDAR sensor used to measure it. That is, the dataset does not provide individual scans. For that reason, using the given scanner poses, we generate per-scanner point clouds by sampling along raycasting paths from each position within the full map, simulating virtual scans. The sampling process respects the scanner’s horizontal angular resolution as well as its vertical resolution between consecutive rays [Ryu23cvpr-InstantDomainAugmentation], ensuring a realistic approximation of the original scans. For each ray, the nearest intersecting point in the merged point cloud is selected to simulate the original scan.
  • ScanNet++i is also from ScanNet++ [Yeshwanth23iccv-Scannet++]. There are depth images captured using a LiDAR sensor attached to an iPhone 13 Pro. These depth images were converted into point clouds using the toolbox provided by 3DMatch [Zeng17cvpr-3dmatch]. To generate dense point cloud fragments, 50 consecutive frames were accumulated. Finally, pairs with an overlap ratio of at least 0.4 (i.e., 40% overlap between two fragments) were selected as the final test pairs.
  • TIERS consists of Indoor06, Indoor08, Indoor09, Indoor10, and Indoor11 sequences in the TIERS dataset [Qingqing22iros-TIERS]. To reduce redundant test pairs, we exclude Indoor07 because it was acquired in the same room as Indoor06.
    Param. Description Value
    $\tau_{v}$ Threshold in Eq. 3 0.05
    $\tau_{l}$ Threshold for the local ($l$) search radius 0.005
    $\tau_{m}$ Threshold for the middle ($m$) search radius 0.02
    $\tau_{g}$ Threshold for the global ($g$) search radius 0.05
    $\delta_{v}$ Sampling ratio for sphericity-based voxelization 10%
    $N_{r}$ Number of sampling points for radius estimation 2,000
    $r_{\text{max}}$ Maximum radius 5.0 m
    $N_{\text{FPS}}$ Number of sampled points by FPS 1,500
    $N_{\text{patch}}$ Maximum number of points in each patch 512
    $\delta$ Truncation threshold for Huber loss 1.0
    $H$ Height of the cylindrical map in Mini-SpinNet 7
    $W$ Sector size of the cylindrical map in Mini-SpinNet 20
    $D$ Feature dimension of the cylindrical map in Mini-SpinNet 32
    1Input: Source cloud $\mathcal{P}$ and target cloud $\mathcal{Q}$; User-defined parameters $\tau_{v}$, $\delta_{v}$, $\delta_{r}$, $\tau_{l},\tau_{m},\tau_{g}]$, and $N_{\text{FPS}}$
@@ 3 $\mathcal{P}_{r}\leftarrow\texttt{select\_larger\_cloud}(\mathcal{P},\mathcal{Q})$
    4 $\mathcal{P}_{\text{sampled}}=\texttt{sample}(\mathcal{P}_{r},\delta_{v})$ % Sample $\delta_{v}$% of cloud points
@@
    6 $v=\texttt{calc\_voxel\_size}(\mathcal{P}_{\text{sampled}},\tau_{v})$ % See Eq. 3
    7 $\mathcal{P}\leftarrow f_{v}(\mathcal{P}),\mathcal{Q}\leftarrow f_{v}(\mathcal{Q})$ % Downsample the point clouds
@@ 8 $\mathcal{P}_{r}\leftarrow\texttt{select\_larger\_cloud}(\mathcal{P},\mathcal{Q})$
    12 $\mathcal{M}^{\mathcal{P}}=\varnothing,\mathcal{M}^{\mathcal{Q}}=\varnothing$ % Containers of embedding output
@@ 23 $\mathcal{D}=\varnothing,\mathcal{T}=\varnothing$
    36 $\mathcal{I}=\texttt{consensus\_maximization}(\mathcal{D},\mathcal{T})$ % See Eq. 9
@@
    Dataset/Sequence Name 3DMatch, 3DLoMatch ScanNet++i ScanNet++F TIERS
@@
    Ouster OS0-128
    Approx. Range [m] 3.5 3.5 7.0 110
    # of Points/Frame 336,274 444,876 1,381,013 34,777
    # of Test Pairs 1,623 / 1,781 2,074 2,016 870
@@
    Follow Zeng et al. [Zeng17cvpr-3dmatch]’s
@@
    Acquisition Vehicle Vehicle Tripod Vehicle Wheeled Quadruped / Handheld
@@
[…]
@@
    Table A2: Overview of the datasets used in our experiments, including their environments, acquisition sites, measurement types, employed sensors, approximate ranges or scales, and the number of test pairs. The datasets cover a variety of indoor and outdoor environments, spanning different geographic and cultural regions. The superscript ✽ highlights solid-state LiDAR sensors to especially emphasize that our evaluation is not limited to conventional omnidirectional spinning LiDAR sensors, but contains different scanning patterns.
  • WOD follows protocol proposed by Liu et al.[Liu24cvpr-Extend]. This dataset is from the Waymo Open Dataset[Sun20cvpr-WaymoDataset] Perception dataset by extracting LiDAR sequences and corresponding pose files, which are then converted into the KITTI format to ensure compatibility with standard benchmarking pipelines. We set $\tau_{\text{dist}}=10\,\text{m}$
  • KITTI follows conventional protocol proposed by Yew et al.[Yew18eccv-3dfeatnet]. In test scenes, 08, 09, and 10 sequences are employed. Originally, $\tau_{\text{dist}}=10\,\text{m}$.
  • ETH is from the gazebo_summer, gazebo_winter, wood_autmn, and wood_summer sequences of the dataset proposed by Pomerleau et al. [Pomerleau12ijrr-ETH]. The original dataset contains a wider various of scenes; however, following the existing protocol proposed by Ao et al. [Ao23CVPR-BUFFER], we used four sequences.
  • KAIST is from the KAIST05 sequence of the HeLiPR dataset [Jung23ijrr-HeLiPR]. Originally, HeLiPR contains multiple sequences, but each sequence in the HeLiPR is much longer than those in MIT and Oxford, resulting in many more test pairs compared to other campus scenes (i.e., 1991 vs. 230 or 301 in Table A2). For this reason, we balanced the datasets by using only one sequence. We set $\tau_{\text{dist}}=10\,\text{m}$.
    Instead of emphasizing that HeLiPR is a heterogeneous LiDAR sensor dataset, we refer to a subset of it as KAIST in our paper to highlight that our dataset was curated with consideration for geographic and cultural environments. Similarly, we use MIT (from the Kimera-Multi dataset [Tian23iros-KimeraMultiExperiments]) and Oxford (from the NewerCollege dataset [Ramezani20iros-NewerCollege]) to emphasize the institutions for the same reason. Further details are provided in Appendix C.
  • MIT is from 10_14_acl_jackal sequence from the Kimera-Multi dataset [Tian23iros-KimeraMultiExperiments], which is a multi-robot multi-session SLAM dataset. We could have used more scenes, but we chose to use only one sequence to a) match Oxford’s frame count as closely as possible (i.e., 230 vs. 301 in Table A2) and b) reduce the redundant test pairs, as multi-robot SLAM datasets often observe the same space multiple times.
@@
  • Oxford is from the 01, 05, and 07 sequences of the NewerCollege dataset [Ramezani20iros-NewerCollege]. An interesting aspect is that 01 and 07 were acquired by handheld setup, while the 05 sequence was acquired using a quadruped robot. As shown in Fig. A2, the campus scale was relatively small, so if we use only a single sequence, it only generates few test pairs. For that reason, to ensure at least a similar number of test pairs as MIT, unlike KAIST and MIT, we used three sequences and set $\tau_{\text{dist}}=5\,\text{m}$.
@@ C Rationale behind dataset selection
Fig. A2: Scale comparison of three sequences: Oxford from the NewerCollege [Ramezani20iros-NewerCollege], MIT from the Kimera-Multi [Tian23iros-KimeraMultiExperiments], and KAIST from the HeLiPR dataset [Jung23ijrr-HeLiPR] at the same scale (i.e., 500 m). Note that although these sequences fall under the same campus category, their scales differ. For clarity, the map clouds are visualized with respect to their $z$ values.
@@
Acquisition setups.   We also considered that acquisition setups vary in multiple ways. A common bias in previous evaluations in the state-of-the-art approaches is that indoor scanning is performed using handheld devices, whereas outdoor scanning is conducted using vehicles (i.e., the assumption that indoor scanning is performed using a handheld setup, while outdoor scanning is conducted using a vehicle). Thus, we aimed to evaluate whether registration remains robust to different acquisition setups to challenge this assumption.
For this reason, we included TIERS, which was acquired using a sensor cart, and Oxford, which was acquired using both a handheld device and a quadruped robot. Additionally, MIT was captured using a mobile robot, which performs planar motion similar to a vehicle. However, due to its smaller size, the robot’s body experiences significantly more roll and pitch motion, potentially introducing greater motion noise compared to a vehicle.
Diversity of geographic and cultural environments. Lastly, we also aimed to assess whether the acquired data could be generalized across geographic and cultural differences. Some studies [Cattaneo22tro-LCDNet] have claimed that training on KITTI and testing on KITTI-360 could demonstrate domain generalization. However, since both datasets were collected in Germany using the same vehicle platform, they fall short of demonstrating actual cultural differences.
@@
Fig. A3: Examples of visualized LiDAR scans from different LiDAR sensors in (a) the Indoor10 of TIERS dataset [Qingqing22iros-TIERS] and (b) KAIST05 sequence of the HeLiPR dataset [Jung23ijrr-HeLiPR]. Note that even in the same environment, differences in the number of LiDAR rays and field of view result in point clouds with different patterns.
@@ D Trade-off between generalization and robustness against partial overlaps
Here, as a supplement to the explanation in Sec. 5.4, we explain with an example why our methodology performed relatively poorly on 3DLoMatch in detail [Huang21cvpr-PREDATORRegistration]. Suppose we have two arbitrary L-shaped point clouds, as presented in Fig. A4(a). Then, these two L-shaped point clouds can be registered in the following cases:
@@
In addition, even with a learnable non-linear robust kernel, this problem cannot be perfectly resolved. For instance, we examine two renowned non-convex cost functions: a) German-McClure (GM) function and b) truncated least squares (TLS) function, and use them as surrogate cost functions $\rho_{\mu}(r)$ that adjust their non-linearity by changing the control parameter $\mu$ [Yang20ral-GNC]. Formally, by letting the $r$ be the residual and $\bar{c}$ be the user-defined parameter that determines the shape of a kernel, the GM function with $\mu$ can be expressed as follows:
@@
Env. Indoor Outdoor Average rank
Conven- tional FPFH [Rusu09icra-fast3Dkeypoints] + FGR [Zhou16eccv-FastGlobalRegistration] + [Uncaptioned image] 62.53 15.42 77.68 92.31 80.60 98.74 100.00 89.80 74.78 91.87 99.00 8.55
FPFH [Rusu09icra-fast3Dkeypoints] + Quatro [Lim22icra-Quatro] + [Uncaptioned image] 8.22 1.74 9.88 97.27 86.57 99.10 100.00 91.46 79.57 51.05 91.03 10.82
FPFH [Rusu09icra-fast3Dkeypoints] + TEASER++ [Yang20tro-teaser] + [Uncaptioned image] 52.00 13.25 66.15 97.22 73.13 98.92 100.00 89.20 71.30 93.69 99.34 8.82
Deep FCGF [Choi19iccv-FCGF] 8.04 0.17 19.96 23.07 77.82 98.92 95.38 88.34 82.17 6.59 75.08 14.82
Predator [Huang21cvpr-PREDATORRegistration] N/A (Err) N/A (Err) N/A (Err) N/A (Err) 69.43 99.82 100.00 N/A (OOM) 54.78 55.68 89.04 14.09
GeoTransformer [Qin23tpami-GeoTransformer] N/A (Err) N/A (Err) N/A (Err) N/A (Err) N/A (Err) 99.82 100.00 63.84 93.91 77.00 73.42 12.73
BUFFER [Ao23CVPR-BUFFER] N/A (Err) N/A (Err) 17.60 88.84 93.34 99.64 100.00 99.50 95.22 98.18 99.34 8.27
PARENet [Yao24iccv-PARENet] N/A (Err) N/A (Err) N/A (Err) N/A (Err) N/A (Err) 99.82 97.69 57.51 75.22 68.30 66.11 15.82
Table A3: Quantitative comparison of generalization performance in terms of success rate (unit: %). Deep learning-based models were trained only on KITTI [Geiger13ijrr-KITTI] and RANSAC was used with a maximum iteration of 50K. Icons represent oracle tuning ([Uncaptioned image]) for voxel size and radius, and scale alignment ([Uncaptioned image]) to match dataset scales (e.g., adjusting 3DMatch’s scale to KITTI by multiplying $\frac{0.3}{0.025}$). N/A (Err) indicates failure due to an insufficient number of points remaining after voxelization with the voxel size typically used for outdoor settings, making keypoint extraction or descriptor generation infeasible. N/A (OOM) indicates an out-of-memory error caused by excessive memory usage.
E Quantitative results using KITTI
Therefore, while training on KITTI led to some variations in overall performance, our approach remains more robust than other state-of-the-art methods.
F Qualitative results in diverse scenes
We present the qualitative results in Fig. A5 and Fig. A6. Remarkably, even when trained on 3DMatch, which consists solely of RGB-D point clouds with an approximate maximum range of 3.5 m, the model performs robustly on sparse LiDAR point clouds in both indoor and outdoor environments. In particular, our approach successfully performed registration regardless of the sensor type or acquisition setup, whether the sparse 3D point clouds were acquired by solid-state LiDAR sensors (the first row among the KAIST rows in Fig. A6) or captured from a robot platform (MIT rows in Fig. A6).